NEED HELP ASA P

Kohler Corporation reports the following components of stockholders’ equity at December 31, 2018. Common stock—$15 par value, 100,000 shares authorized, 45,000 shares issued and outstanding $ 675,000

Paid-in capital in excess of par value, common stock 70,000

Retained earnings 430,000

Total stockholders' equity $ 1,175,000

During 2019, the following transactions affected its stockholders’ equity accounts.

Jan. 2 Purchased 6,000 shares of its own stock at $15 cash per share.

Jan. 5 Directors declared a $2 per share cash dividend payable on February 28 to the February 5 stockholders of record.

Feb. 28 Paid the dividend declared on January 5.

July 6 Sold 2,250 of its treasury shares at $19 cash per share.

Aug. 22 Sold 3,750 of its treasury shares at $12 cash per share.

Sept. 5 Directors declared a $2 per share cash dividend payable on October 28 to the September 25 stockholders of record.

Oct. 28 Paid the dividend declared on September 5.

Dec. 31 Closed the $408,000 credit balance (from net income) in the Income Summary account to Retained Earnings.

Required:

1. Prepare journal entries to record each of these transactions.

2. Prepare a statement of retained earnings for the year ended December 31, 2019.

3. Prepare the stockholders’ equity section of the company’s balance sheet as of December 31, 2019.
Answer #1
1
Date General Journal Debit Credit
2-Jan Treasury stock, Common 90,000 =6000*15
Cash 90,000
5-Jan Retained earnings 78,000 =(45000-6000)*2
Common dividend payable 78,000
28-Feb Common dividend payable 78,000
Cash 78,000
6-Jul Cash 42,750 =2250*19
Treasury stock, Common 33,750 =2250*15
Paid-in capital, Treasury stock 9,000
22-Aug Cash 45,000 =3750*12
Paid-in capital, Treasury stock 9,000
Retained earnings 2,250
Treasury stock, Common 56,250 =3750*15
5-Sep Retained earnings 90,000 =45000*2
Common dividend payable 90,000
28-Oct Common dividend payable 90,000
Cash 90,000
31-Dec Income summary 408,000
Retained earnings 408,000
2
KOHLER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2019
Retained earnings, Dec. 31, 2018 430,000
Add: Net income 408,000
838,000
Less: Cash dividends declared -168,000
Less: Treasury stock reissuances -2,250
Retained earnings, Dec. 31, 2019 667,750
3
KOHLER CORPORATION
Stockholders' Equity Section of the Balance Sheet
December 31, 2019
Common stock - $15 par value 675,000
Paid-in capital in excess of par value, common stock 70,000
Total contributed capital 745,000
Retained earnings 667,750
Total stockholders' equity 1,412,750